# Subject-Oriented Video Captioning

### v1

@@ Abstract
Describing video content according to users’ needs is a long-held goal. Although existing video captioning methods have made significant progress, the generated captions may not focus on the entity that users are particularly interested in. To address this problem, we propose a new video captioning task, subject-oriented video captioning, which allows users to specify the describing target via a bounding box. To support this task, we construct two subject-oriented video captioning datasets based on two widely used video captioning datasets: MSVD and MSRVTT, by annotating subjects in each video for each caption. These datasets pave the way for future technique development. As the first attempt, we evaluate four state-of-the-art general video captioning models, and have observed a large performance drop. We then explore several strategies to enable them to describe the desired target. Experimental results show obvious improvement, but there is still a large room for further exploration in this field.
1 Introduction
@@
Video captioning is an important video understanding task, aiming to accurately describe the video content in a single sentence. It plays a crucial role in numerous applications, such as video title generation [35], blind assistance [6], and video search [13]. However, existing methods do not allow users to specify the target of interest so as to generate desired captions. This could largely limit its application in real-world scenarios. Consider a video where multiple individuals are present as shown in Figure 1, if a user needs a summary of the behaviors of a specific individual, such as the woman, traditional video captioning methods have a large chance to fail, e.g., the state-of-the-art methods SwinBERT [9] and HMN [32] generate “A man is talking about a car” and “A man is driving a car”, respectively. This is because they are trained to cover the majority of entities mentioned in the ground truth, which is “a man” in this case. Consequently, they cannot cater to user-specific needs for targeted descriptions.
@@
In the field of image captioning, similar attempts have been made. For instance, a framework [4] named “Show, Control and Tel” generates captions by taking a series of image regions as input. However, this method requires users to select and rank multiple entity areas, which is impractical in the video scenario because it is unreasonable to require the user to watch the whole video first and then pick out interacted objects. A desired method should just require to specify the target of interest and let the model its own to discover what happens to the target. Another limitation of applying image captioning methods for video captioning is their inability to handle temporal information in multiple frames, such as complex actions. The above mentioned limitations also exist in other image captioning methods, such as CGO [37] and VSR-guided model [3].
@@
Another line of related research is controlled video captioning. Existing methods primarily focus on controlling sentence structure, as demonstrated by [25], which introduces Part-of-Speech(POS) information for this purpose. SCVC [34] aims to generate one caption that describes the video contents semantically while also imitating the syntactic structure of the given example sentence. However, these methods lack effective control over the subjects in the videos. Meanwhile, although O2NA [11] allows for control using simple entity-class words, it does not enable the selection of specific entity areas, thus lacking flexibility and failing to meet our requirements.
@@
To address the aforementioned problems, we propose a new task named “subject-oriented video captioning”, which aims to generate descriptions centered on activities of the user-specified entity. One key reason limiting the development of such desired methods is the lack of proper data. Existing video captioning datasets do not provide video annotations corresponding to the subject in each caption. This raises difficulty in learning the correspondence between captions and numerous entities in a video. In this work, we annotate two widely used video captioning datasets: MSVD [2] and MSR-VTT [26]. This paves the way to develop methods that generate captions describing user-specified targets. As an example, in the bottom panel of Figure 1, trained with the new data, our method can generate descriptions closely related to each specified entity. For example, when specifying the car as the description target, the generated caption is “A car is being shown”; when the target of interest is changed to the woman, the generated caption changes to “A woman is driving a car”.
To verify the necessity of developing new methods in this direction, we evaluate four state-of-the-art conventional video captioning methods and observe significant performance degradation on the new task: 20$\sim$30 absolute drop in terms of the main metric Cider on our subject-oriented MSVD, and 6$\sim$51 absolute drop on our subject-oriented MSRVTT. This shows there is ample room for exploration in this field. As our initial endeavor, we have undertaken several preliminary attempts where we integrated subject features into existing frameworks based on Transformer and LSTM. These preliminary attempts have yielded certain successes and may serve as inspiration for future research.
The main contributions of this paper are as follows:
  • We propose a new task, subject-oriented video captioning, and build two new datasets based on MSVD and MSR-VTT, which paves the way for developing more flexible video captioning methods.
  • We evaluate four representative conventional video captioning methods, and observe large performance degradation, which shows the necessity of developing new techniques.
  • We carry out extensive preliminary attempts to address the new task and achieve favorable improvements, which may serve as a strong baseline and inspire future research.
@@
2 Related Work
In this section, we briefly review general video captioning methods and controllable video captioning methods, which are related to our task and put this work in a proper context.
2.1 General Video Captioning
Recent research on video captioning primarily focuses on the design of network architectures. Early methods employ simple CNN architectures for representation learning from videos, such as S2VT [24]. Later, more complex encoder architectures are proposed. STG-KD [14] and ORG-TRL [36] design encoders based on object relation graphs to distinguish the spatial-temporal relationships among video objects. SGN [19] proposes a semantic grouping network, which encodes videos into semantic clusters based on relevant frames. HMN [32] adopts a hierarchical modular design to encode three levels of visual features to get comprehensive video representations. GL-RG [27] creates a global-local encoder, which exploits rich temporal representation for video captioning. Some work also focuses on improving the decoding ability, such as the integration of temporal attention mechanisms by Yap et al. [31] and attended memory decoders by Pei et al. [16], marking significant enhancements in captioning quality. In addition to obtaining richer visual representations and enhancing decoding capabilities, the long-tailed problem hinders the effectiveness of low-frequency tokens, which play a crucial role in generating detailed content. RSFD [38] is proposed to constantly capture the linguistic representation of the infrequent tokens. TextKG [5] integrates an additional knowledge graph to mitigate the challenge of long-tail words.
Different from the aforementioned methods, SwinBERT [9] employs an end-to-end transformer-based design, which directly takes a sequence of video frames as inputs and uses caption-related loss to adjust feature learning. However, it increases the computational burden and time consuming to handle the dense video tokens. To enhance inference efficiency, Cocap [21] directly takes compressed video as input and fast generates captions. Most recently, some work [7, 20] leverages large pre-train models as the backbone to align powerful vision and language representations.
@@ 2.1 General Video Captioning
2.2 Controllable Video Captioning
In contrast to general video captioning, much less attention is paid to controllable video captioning. Some efforts involve using additional sentences as reference signals, allowing the generated sentences to learn the syntactic structure or style of the reference sentences. For example, SMCG [33] models the syntactic structure from a given exemplar sentence and designs a modified LSTM module to transfer the syntax in the exemplar to new captions for the input video. FS-StyleCap [29] extract the style of stylized example to generate sentences with the same style. On the other hand, some works employ special inputs as control signals. MSG [30] uses a controllable masked scene graph to represent interested objects and develops a corresponding decoder to infer masked relationships between objects to obtain the final caption. In [11], a method to generate captions about the specified objects is developed. To this end, object detection is performed, and then a non-autoregressive generator has to be used to make up a complete sentence from ahead-determined objects. ERA [10] introduces the reference videos to capture contextual information and further refine the group video captioning.
@@ 2.2 Controllable Video Captioning
3 Subject-Oriented Video Captioning Task
In this paper, we introduce a new task, termed as “Subject-Oriented Video Captioning”. Unlike conventional video captioning tasks that generate a broad description of a video, our task aims to generate captions that predominantly focus on the activities of a specific entity specified by the user using a bounding box. To adapt to this newly devised task, traditional video captioning datasets have been meticulously refined, annotated, and reorganized, ensuring that they are well-suited to facilitate and nurture the generation of subject-oriented video captions. This enhances the specificity and relevance of the textual descriptions, echoing the nuanced activities of the identified subjects in videos.
In the next section, we provide the details of the dataset construction and statistics. At last, we evaluate four state-of-the-art conventional video captioning methods and our extensions on top of them.
@@ 3 Subject-Oriented Video Captioning Task
Table 1: Statistics of our Subject-Oriented MSVD and MSRVTT datasets.
4 Dataset
4.1 Dataset Construction
To facilitate future research on the Subject-Oriented Video Captioning task, we construct datasets based on two widely used general video captioning datasets: MSVD [2] and MSRVTT [26].
@@ 4.1 Dataset Construction
4.2 Dataset Statistics
Table 1 presents the detailed specifics of the reannotated and reorganized datasets. In the subsequent sections, we elaborate on each of them, drawing upon the data shown in the table.
Subject-Oriented MSVD After filtering out captions with abstract subjects from the original MSVD dataset [2], we annotate 3,308 subject samples for 1,961 videos, with 72,530 captions. On average, one subject corresponds to 22 sentences. As shown in Figure 3, subject-oriented MSVD contains a varying number of subjects ranging from 1 to 9, with the corresponding average number of captions approximately between 3 and 38. Note that one subject may correspond to multiple frames and regions in a video. For example, consider the scenario where two men are talking about something as shown in Figure 4; these ‘two men’ may first appear in different frames, and each may occupy different regions even in the same frame. Based on our statistics, a total of 4,287 regions have been annotated across 3,411 frames. Since the total number of videos remains similar to its original scale, we adopt the same dataset division, leading to 1871, 150, and 1287 subject samples in the training, validation, and test splits, respectively. We also provide word-cloud-based statistics in Figure 5 (a) to reveal the relative amount of different subjects. It shows that the top-4 subjects in MSVD are “man”, “girl”, “person”, and “cat”, followed by “boy”, “dog”, “men”, and “monkey”.
@@ 4.2 Dataset Statistics
Subject-Oriented MSRVTT We apply the same annotation strategy to the newly built training, validation, and test splits contain 15,306, 1,085, and 6,979 subject samples, respectively, from 6,310, 487, and 2,927 videos, respectively. Figure 3 illustrates the range of subject numbers in each video of subject-oriented MSRVTT, varying from 1 to 10, with the corresponding average number of captions approximately between 1 and 15. On this dataset, a total of 35,499 regions have been annotated across 25,988 frames. Figure 5 (b) presents the relative amount of different subjects in the reconstructed MSRVTT dataset. It shows that the top-4 subjects are slightly different from those of the new MSVD, where ‘people’ substitutes ‘cat’. We also observe that ‘men’, ‘chef’, ‘car’, ’player’, ’guy’, ‘women’, ‘children’, and ‘character’ have a similar amount, just next to the top-4 subjects.
(a) Subject-Oriented MSVD
(b) Subject-Oriented MSRVTT
@@
5 Method
Figure 6: The top row shows the original method, while the bottom row displays our improved approach.
To test the difficulty of the subject-oriented video captioning task, we choose four representative state-of-the-art video captioning methods, including AR-B [28], HMN [32], GL-RG [27] and SwinBERT [9]. According to the decoder architecture, HMN and GL-RG both adopt LSTM-based frameworks, whereas AR-B and SwinBERT are representative approaches utilizing Transformers [22]. Moreover, we also extended these four baselines to make them more suitable to address the newly proposed task.
In the following sections, we first briefly introduce the four baselines in Sec. 5.1. Then, we describe the specific adaptation of LSTM-decoder-based models to our task in Sec 5.2. Last, we present the modification of transformer-based methods in Section 5.3.
5.1 Baseline Methods

### v2

# SOVC: Subject-Oriented Video Captioning

@@ Abstract
Describing video content according to users’ needs is a long-held goal. Although existing video captioning methods have made significant progress, the generated captions may not focus on the entity that users are particularly interested in. To address this problem, we propose a new video captioning task, Subject-Oriented Video Captioning (SOVC), which aims to allow users to specify the describing target via a bounding box. To support this task, we construct two subject-oriented video captioning datasets based on two widely used video captioning datasets: MSVD and MSRVTT, by annotating subjects in each video for each caption. These datasets pave the way for describing users’ interested targets. To tackle this task, we introduce a method tailored to this task, named SOVCNet. It consists of two key components: a subject-oriented sampling module that samples frames related to the subject to minimize irrelevant information; and a subject-oriented encoding module that utilizes the subject areas as hard prompts and integrates learnable soft prompts, enhancing the model’s focus on the subject’s activities and facilitating adaptation to the downstream generation task. Extensive experimental results demonstrate the effectiveness of our method on this new task.
Index Terms:
Subject-oriented, Video captioning, Prompt learning
I Introduction
Video captioning [1, 2, 3, 4, 5, 6, 7] is an important video understanding task, aiming to understand and accurately describe the video content in a single sentence. It plays a crucial role in numerous applications, such as video title generation [8], blind assistance [9], and video search [10]. However, existing methods do not allow users to specify the target of interest and therefore the generated captions might not be the desired. This could limit its application in real-world scenarios. Consider a video where multiple individuals are present, as shown in Figure 1,
@@ I Introduction
if a user needs a summary of the behaviors of a specific individual, such as the woman, traditional video captioning methods have a large chance to fail, the state-of-the-art methods SwinBERT [11] and HMN [12] generate “A man is talking about a car” and “A man is driving a car”, respectively. This is because they are trained to cover the majority of entities mentioned in the ground truth, which is “a man” in this case. Consequently, they cannot cater to user-specific needs for targeted descriptions.
@@
In the field of image captioning [13, 14, 15, 16, 17], similar attempts have been made. For instance, a framework [18] named “Show, Control and Tell” generates captions by taking a series of image regions as input. However, this method requires users to select and rank multiple entity areas, which is less practical in the video scenario because it is unreasonable to require the user to watch the whole video first and then pick out interacted objects. A desired method should just require to specify the target of interest and let the model its own to discover what happens to the target. Another limitation of applying image captioning methods for video captioning is their inability to handle temporal information in multiple frames, such as complex actions. The above-mentioned limitations also apply to other image captioning methods, such as CGO [19] and VSR-guided model [20].
@@
Another line of related research is controlled video captioning. Existing methods primarily focus on controlling sentence structure, as demonstrated by [21], which introduces Part-of-Speech (POS) information for this purpose. SCVC [22] aims to generate one caption that describes the video contents semantically while also imitating the syntactic structure of the given example sentence. However, these methods lack effective control over the subjects in the videos. Meanwhile, although O2NA [23] allows for control using simple entity-class words, it lacks the ability to select a specific entity if multiple instances of the same class exist, thus lacking specification flexibility and precision.
@@
To address the aforementioned problems, we propose a new task named “subject-oriented video captioning”, which aims to generate descriptions centered on activities of the user-specified entity. As an example, in the bottom panel of Figure 1, when specifying the car as the description target, the generated caption is “A car is being shown”; when the target of interest is changed to the woman, the generated caption changes to “A woman is driving a car”. However, compared to conventional video captioning, the new task faces two main challenges: 1) how to capture the most relevant information related to the subject in the video. 2) how to model the relationship between the subject and video information.
To tackle these challenges, we propose a subject-oriented video captioning Network (SOVCNet), which incorporates two of our proposed modules (subject-oriented sampling module and subject-oriented encoding module) into existing methods. Conventional video captioning methods uniformly sample fixed video frames to perceive the overall video content. In our task, we argue the effectiveness of traditional sampling methods because our objective is to describe the subject’s activities rather than the overall video content. Thus, we develop a new subject-oriented sampling method specifically designed for our proposed task. We first use an image encoder to extract features of the subject and frames and compute their cosine similarity. Then, based on the cosine similarity, we use a cluster-based method to sample a fixed number of frames. After sampling subject-related video frames, we further consider how to leverage these visual information to model the relationship between the subject and other entities in the video. Inspired by the success of P-Tuning [24, 25] and Visual Prompt Tuning (VPT) [26], we design a new subject-oriented encoding module. This module utilizes subject features as hard prompts to direct the video encoder’s attention toward the subject’s activities. Additionally, it plugs learnable soft prompts that enable the model to adapt to the generation task.
To support our proposed task, we re-annotate two widely used video captioning datasets: MSVD and MSR-VTT. Specifically, we identify frames that match the video description and label the coordinates of the subjects. More details about the annotations and dataset are provided in Section III.
Our main contributions can be summarized as:
  1. 1.
    We propose a novel task: subject-oriented video captioning, which allows the users to focus on objects of interest, especially in some real-world application scenarios, such as surveillance videos, etc.
  2. 2.
    We propose a new model for subject-oriented video captioning, which contains two modules: the subject-oriented sampling module and the subject-oriented encoding module. The former samples frames related to the subject to reduce unnecessary information. The latter uses the subject area as hard prompts and plugs the soft prompts to make the model focus more on the subject’s activities and adapt to the generation task.
  3. 3.
    We constructed two new datasets named SO-MSVD and SO-MSRVTT specifically for the subject-oriented video captioning problem.
  4. 4.
    Extensive experiments on the newly built datasets demonstrate our model outperforms various baselines for the subject-oriented video captioning task.
@@
II Related Work
In this section, we briefly review general video captioning, controllable video captioning and prompt learning, which are related to our task and put this work in a proper context.
II-A General Video Captioning
Recent research on video captioning primarily focuses on the design of network architectures. Early methods apply simple CNN architectures for representation learning from videos, such as S2VT [27]. Later, more complex encoder architectures are proposed. STG-KD [28] and ORG-TRL [29] develop encoders using object relation graphs to distinguish the spatial-temporal relationships among video objects. SGN [30] proposes a semantic grouping network, which encodes videos into semantic clusters by identifying relevant frames. HMN [12] adopts a hierarchical modular design to encode three levels of visual features to get comprehensive video representations. GL-RG [31] creates a global-local encoder, which exploits rich temporal representation for video captioning. Meanwhile, some work also focuses on improving the decoding ability, such as Yap et al. [32] integrating temporal attention mechanisms and Pei et al. [33] incorporating attended memory decoders. Beyond obtaining richer visual representations and enhancing decoding capabilities, the long-tailed problem remains a challenge, as low-frequency tokens are crucial for generating detailed content. RSFD [34] addresses this by persistently capturing the linguistic representations of these rare tokens, while TextKG [35] integrates an additional knowledge graph to mitigate the challenge of long-tail words.
Different from the above methods, SwinBERT [11] employs an end-to-end transformer-based design, which directly takes a sequence of video frames as inputs and uses caption-related loss to adjust feature learning. However, it increases the computational burden and time consuming to handle the dense video tokens. To improve inference efficiency, Cocap [36] directly uses compressed video as input, enabling faster caption generation. Most recently, some work [37, 38] leverages large pre-train models as the backbone to align powerful vision and language representations.
@@ II-A General Video Captioning
II-B Controllable Video Captioning
In contrast to general video captioning, much less attention is paid to controllable video captioning. Some efforts involve using additional sentences as reference signals, allowing the generated sentences to learn the syntactic structure or style of the reference sentences. For example, SMCG [39] models the syntactic structure from a given exemplar sentence and designs a modified LSTM module to transfer the syntax in the exemplar to new captions for the input video. FS-StyleCap [40] extract the style of stylized example to generate sentences with the same style. SAVC [41] proposes a two-stage framework to learn seven different styles and generate more specific captions. On the other hand, some works employ special inputs as control signals. MSG [42] uses a controllable masked scene graph to represent interested objects and develops a corresponding decoder to infer masked relationships between objects to obtain the final caption. In [23], a method to generate captions about the specified objects is developed. To this end, object detection is performed, and then a non-autoregressive generator has to be used to make up a complete sentence from ahead-determined objects. ERA [43] introduces the reference videos to capture contextual information and further refine the group video captioning.
@@ II-B Controllable Video Captioning
II-C Prompt Learning
In recent years, prompt learning has made significant progress in the NLP field. Prompts can be separated into hard prompts and soft prompts. The former usually uses fixed text templates to guide the model in producing the desired responses, which are direct and clear. GENPET [44] and PET [45] employ pre-defined templates in a few-shot learning framework for text classification and conditional text generation tasks. Wallace et al. [46] applies a gradient-based search over actual tokens to automatically find the suitable templates described in a discrete space. The latter employs learnable parameters to refine the model’s adaptation to specific downstream tasks. Prefix-tuning [47] delivers performance comparable to full fine-tuning while reducing parameter storage by a factor of 1000. Additionally, prompt-relevant parameters can also be fine-tuned together with all or some of the parameters of the pre-trained models. PADA [48] and P-Tuning [24] are notable examples. This method can better adapt to downstream tasks and achieve optimal performance.
Beyond the NLP domain, prompt learning also demonstrates its effectiveness in computer vision tasks. For example, Visual Prompt Learning [26] utilizes visual prompts to enhance pre-trained vision models, enabling specific tasks to be performed more effectively with less training data. Besides, ProTrack [49] facilitates multimodal tracking using non-learnable prompts.
Inspired by the aforementioned work, we integrate subject information as hard prompts and learnable soft prompts to enhance the model’s understanding of the user-specified subject and adapt to the downstream generation task.
III Dataset
To support our proposed subject-oriented video captioning task, we require a corresponding dataset that meets our research needs. Each video in the dataset includes $N$ chosen subjects, with corresponding activity descriptions for each subject.
TABLE I: Statistics of our Subject-Oriented MSVD and MSRVTT datasets.
@@ III Dataset
III-A Dataset Construction
We first construct two datasets based on two widely used general video captioning datasets: MSVD [50] and MSRVTT [51].
@@ III-A Dataset Construction
III-B Dataset Statistics
Table I presents the detailed specifics of the reannotated and reorganized datasets. In the subsequent sections, we elaborate on each of them, drawing upon the data shown in the table.
@@ III-B Dataset Statistics
Subject-Oriented MSVD After filtering out captions with abstract subjects from the original MSVD dataset [50], we annotate 3,308 subject samples for 1,961 videos, with 72,530 captions. On average, one subject corresponds to 22 sentences. As shown in Figure 3, subject-oriented MSVD contains a varying number of subjects ranging from 1 to 9, with the corresponding average number of captions approximately between 3 and 38. Note that one subject may correspond to multiple frames and regions in a video. For example, consider the scenario where two men are talking about something as shown in Figure 4; these ‘two men’ may first appear in different frames, and each may occupy different regions even in the same frame. Based on our statistics, a total of 4,287 regions have been annotated across 3,411 frames. Since the total number of videos remains similar to its original scale, we adopt the same dataset division, leading to 1871, 150, and 1287 subject samples in the training, validation, and test splits, respectively.
@@
Subject-Oriented MSRVTT We apply the same annotation strategy to the newly built training, validation, and test splits contain 15,306, 1,085, and 6,979 subject samples, respectively, from 6,310, 487, and 2,927 videos, respectively. Figure 3 illustrates the range of subject numbers in each video of subject-oriented MSRVTT, varying from 1 to 10, with the corresponding average number of captions approximately between 1 and 15. On this dataset, a total of 35,499 regions have been annotated across 25,988 frames. Figure 5(b) presents the relative amount of different subjects in the reconstructed MSRVTT dataset. It shows that the top-4 subjects are slightly different from those of the new MSVD, where ‘people’ substitutes ‘cat’. We also observe that ‘men’, ‘chef’, ‘car’, ‘player’, ‘guy’, ‘women’, ‘children’, and ‘character’ have a similar amount, just next to the top-4 subjects.
@@
Figure 6: An overview of our framework. For the video input, we first perform subject-oriented sampling (SOS) to sample the frames related to the subject. Then, we leverage the subject-oriented encoding (SOE) to enhance the model’s focus on the subject’s activities, tailoring it more effectively to the generation task. Finally, the encoded tokens and subject tokens are concatenated and then fed into the caption generator for description generation.
IV Method
In this section, we first formalize our proposed SOVC task (Sec. IV-A). We then simply extend the existing methods to our task as the baseline (Sec. IV-B). Finally, we introduce our proposed SOVCNet, which mainly contains the subject-oriented sampling module and subject-oriented encoding module (Sec. IV-C)
IV-A Task definition
Given a test video $\mathbf{V}=[\mathbf{I}_{1},\ldots,\mathbf{I}_{M}]$ of $M$ frames, traditional video captioning methods uniformly sample fixed $N$ frames to represent the whole video and learns a model to describe it. Differently, in this paper, our purpose is to select areas of interest in images as subjects and to learn a new model to describe their activities within the video.
IV-B Simplifying Extension of Existing Methods for Our Task
The workflow of traditional video captioning methods can be described as follows:
$\mathbf{C_{\mathbf{V}}}=G_{lan.}(E_{\mathbf{V}}(\mathbf{V}))$ (1)
where $E_{\mathbf{V}}$ and $G_{lan.}$ represent the video encoder and language generator, $\mathbf{C_{\mathbf{V}}}$ is the generated caption about $\mathbf{V}$. To adapt existing methods to our task, a straightforward idea is to merge subject features with video features, using this combination as the input for the decoder to generate sentences related to the subject. The entire process is illustrated as follows:
$\mathbf{C_{\mathbf{S}}}=G_{lan.}(E_{\mathbf{V}}(\mathbf{V}),E_{\mathbf{S}}(\mathbf{S}))$ (2)
where $\mathbf{S}$ and $E_{\mathbf{S}}$ are subject area and subject encoder, and $\mathbf{C_{\mathbf{S}}}$ is the generated caption describing the activities of subject.
However, this approach faces two main challenges: 1) the original uniform video frame sampling method tends to incorporate information irrelevant to the subject, and 2) the modeling of relationships between the subject and video frames is overly direct, making it difficult to adapt to downstream generation tasks. Therefore, we use this straightforward improvement as a simple baseline and introduce the modules for our further enhancements.
IV-C SOVCNet
We propose a new method named SOVCNet, which first samples video frames related to the selected subject (Section. IV-C1), then encodes the input obtained from the previous step (Section. IV-C2), finally generates sentences that describe the activities of the subject (Section. IV-C3).
IV-C1 Subject-Oriented Sampling
Conventional sampling methods capture one frame at regular intervals, such as every 8 or 32 frames. However, in the SOVC task, frames that contain information correlated to the subject need to be exploited. To preserve diverse and representative subject-related information, we proposed a subject-oriented sampling strategy, which contains two steps, i.e. clustering and sampling.
Specifically, we first extract features of frames using Resnet50 [52] because of its high efficiency. Let $f_{\text{subject}}$ and $D=\{f_{1},f_{2},\ldots,f_{N}\}$ denote the features of the subject frame and the set of all frames in the video, respectively. Then, we adopt K-means [53] to cluster $D$ into $T$ clusters $C=\{C_{1},C_{2},\ldots,C_{T}\}$, where $T$ is the number of frames that need to be selected.
To ensure the diversity of sampled video frames, we sample subject-related frames in each cluster. For each frame, its cosine similarity to the subject frame is calculated as,
$s(f_{i})=\text{sim}(f_{i},f_{subject}),i=1,2,\cdots,N$ (3)
Convert these similarities to probabilities using the softmax function:
$p(f_{i}\mid C_{j})=\frac{e^{s(f_{i})}}{\sum_{f_{k}\in C_{j}}e^{s(f_{k})}},\quad\forall f_{i}\in C_{j},\text{ for each }C_{j}\in C$ (4)
Then, from each cluster, We sample one frame based on probability. For simplicity, we use $F\in\mathbb{R}^{T\times H\times W\times C}$ to represent the $T$ sampled frames.
IV-C2 Subject-Oriented Encoding
In the encoding phase, we propose a subject-oriented encoding module to make the model focus more on the subject’s activities and adapt to the generation task. As shown in Figure 7 (a), the conventional encoding module directly uses the patch embedding operation to process the video frames into a sequence of tokens. In contrast, we design a subject-oriented encoding module to enable the video encoder to understand the subject we are focusing on. Specifically, as shown in Figure 7 (b), in addition to performing patch embedding on video frames to obtain a series of tokens, we apply the same operation to the subject area and use it as a visual hard prompt to guide the model’s focus on the subject’s activities. Meanwhile, we add some learnable tokens as soft prompts to better adapt the model for downstream generative tasks. Then, we feed the joint representation of these tokens into the video encoder to obtain a more appropriate feature representation. The entire process can be described as follows:
$\bar{\mathbf{V}}=E_{\mathbf{V}}(Concat(PE(\mathbf{V},\mathbf{S}),\mathbf{P_{soft}}))$ (5)
where $PE$ denotes the patch embedding operation, $\mathbf{P_{soft}}$ is the learnable soft prompt, $\bar{\mathbf{V}}$ is the encoded video feature representation.
Figure 7: Comparison between the Conventional Encoding (a) and Our Subject-Oriented Encoding (b).
By combining SOS and SOE modules, our model is capable of fully perceiving the information in the video that is most relevant to the subject, thereby generating more accurate descriptions of the subject.
IV-C3 Subject-Oriented Caption Generation
To generate the subject-oriented caption, we incorporate additional subject information into the caption generation module. As shown in Figure 6, we use a pre-trained feature extractor and a linear layer as the subject encoder to extract features of the selected subject area and adjust dimensions. Then, the subject features and video features are concatenated as the input of the language generator to generate the subject-oriented caption $C_{\mathbf{S}}$:
$C_{\mathbf{S}}=G_{lan.}(Concat(\bar{\mathbf{V}},E_{\mathbf{S}}(\mathbf{S})))$ (6)
This step ensures that the generated caption are aligned with the user-specific subjects in the video.
V Experiments
V-A Experimental settings
V-A1 Metrics
V-A2 Baseline models
As described in Sec. IV-B, we directly extend four existing state-of-the-art methods to our task. The descriptions of the original four methods are as follows: